Parallel to the line 4x + 10y = 20 and passes through the point (1,0).
Answer by AnlytcPhil(1806)

Equations of parallel lines in standard form Ax+By=C only need differ by the
constant C.  Thus the line through 

(1,0)

parallel to the line with equation

4x + 10y = 20

can have an equation with the same left side, 4x + 10.

All we need do is substitute (1,0) in the left side

4(1) + 10(0) = 4 + 0 = 4

to find the constant.

So one answer could be

4x + 10y = 4

However since that can be divided through by 2, we should
do so and get

2x + 5y = 2      <--answer

Edwin